# Enhancing Large Language Models for Text-to-Testcase Generation

### v1

@@ 5 Experimental Results
In this section, we present the results with respect to our four research questions.
RQ1: How effective is the proposed approach for the Text-to-Test generation task?
Figure 5: The experimental results of the ablation study with six different models
@@
Results. Our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) outperforms all other LLMs (basic GPT-3.5-turbo, Bloom, and CodeT5). Figure 5 presents the results of our approach and the baseline models for a text-to-testcase generation task. We find that our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) is able to generate test cases that are 78.5% being syntactically correct, 67.09% being aligned with the requirements while achieving a code coverage of 61.7%, which substantially outperforms all other LLMs (basic GPT-3.5-turbo, Bloom and CodeT5). These results highlight the significant advancement of our approach for the text-to-testcase generation task when compared to other baseline LLMs.
@@
The basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task. For CodeT5 and Bloom with the other three scenarios (i.e., No Fine-Tune +Basic P, No Fine-Tune + Imp P, Fine-Tune + Basic P), these models achieve a very low-performing performance (e.g., less than 1% syntax correctness of the generated test cases), meaning that they are not capable to be used in practices. When comparing the BloomFineTune+Imprv.P model and the GPT3.5 NoFineTune+Basic.P model, we found that the basic usage of ChatGPT (GPT-3.5-turbo with no fine-tuning and a basic prompt) may not achieve a desirable performance as it performs worse than the fined-tuned Bloom with the improved prompt design.
Nevertheless, our approach (the GPT3.5 FineTune+Imprv.P model) still performs the best.
RQ2: What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
@@
Results. When fine-tuned the LLMs (GPT-3.5-turbo) with our curated fine-tuned dataset, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT. When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see Figure 5). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 41.60% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
@@
RQ3: What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
@@
Results. With our improved prompts, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT. When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see Figure 5). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 28.90% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of requirement alignment, we found that the performance is improved from 11.10% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 20.25% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of code coverage, we found that the performance is improved from 10.44% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 16.43% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
These results confirm that the improved prompt substantially improves the performance of GPT-3.5, suggesting that prompting design is a crucial step when using a large language model like GPT-3.5, which should be considered in the future.
@@
RQ4: What types of errors occur in the incorrectly generated test cases?
@@
Results. The majority of the generated test cases by our approach are syntactically correct for 78.5%, while being incorrect for 21.5%. The most common types of errors are assertion errors (11.3%), value errors (2.4%), syntax errors (6.9%), and others (0.9%). Understanding the incorrectly generated test cases by our approach is crucial as it helps researchers gain a better understanding of the limitations of our approach. To do so, we utilized Apache Maven (maven, 2023) as our build tool, and the Surefire Plugin (surefire, 2023) to execute the test cases and to analyze the types of errors during the build process. Figure 6 provides a statistical summary of the error types of the generated test cases. An example of different types of incorrectly generated test cases is also presented in Listing 1. Below, we discuss the different types of errors.
@@
  • 1.
  • 2.
  • 3.
[…]
  • 4.
    The other (0.9%) category probably includes all other types of errors that do not fit into the aforementioned categories.
@@ 6 Threats to Validity
In our fine-tuning step, we employed a variety of hyperparameter configurations, including adjustments in the number of epochs, batch size, and learning rate. The hyperparameter optimization can significantly improve evaluation outcomes (Bischl et al., 2023).
However, identifying the optimal settings is resource-intensive due to the expansive search space of hyperparameters. We note that our primary objective in this study is not to pinpoint the most effective hyperparameters but to investigate the potential of LLMs in TDD. Thus, while the results achieved with LLMs in our study are promising, they only represent a starting point and have the potential for further improvement through extensive hyperparameter tuning. Nonetheless, we have included detailed information about the hyperparameter settings used in our study within our replication package. The second major concern related to LLMs is the training data leakage (Xia et al., 2023). There is a possibility that GPT-3.5-turbo or other LLMs may have been exposed to the same open-source projects that we used in our study. However, the pre-trained models did not yield satisfactory results, compared to the fine-tuned model and custom prompt results. This shows that the impact of training data leakage, if any, was minimal in our study settings.
Construct Validity. Our study’s main construct validity concern is related to the prompt tuning and robustness (Lester et al., 2021). Given the large scope of subjectivity in prompting and the lack of empirical guidelines on prompt tuning and robustness in SE, one can always further tune and build a more robust prompt to achieve better evaluation results. However, we experimented with several prompt combinations in our journey, from the basic to the custom prompt, in an attempt to mitigate the threats related to prompt tuning.

### v2

@@ 5 Experimental Results
In this section, we present the results with respect to our three research questions.
RQ1: How effective is the proposed approach for the Text-to-Testcase generation task?
@@
Results. Our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) outperforms all other LLMs (basic GPT-3.5-turbo, Incoder, Starcoder, CodeT5, Bloom, CodeGemma, CodeLlama, and Gemini. Table 4 presents the results of our approach and the baseline models for a text-to-testcase generation task.
The application of fine-tuning and prompting methodologies did not consistently enhance the performance of various Large Language Models (LLMs) such as Incoder, Starcoder, CodeT5, these models achieve a very low-performing performance (e.g., less than 1-10% syntax correctness of the generated test cases), meaning that they are not capable to be used in practices in the domain of test case generation using TDD. The other models such as Bloom, CodeGemma, CodeLama, Gemini, and GPT-3.5 models exhibit diverse efficacy across multiple evaluation metrics including syntax correctness, requirement alignment, code coverage, and mutation scores.
@@
The basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task. When comparing the GPT3.5FineTune+Imprv.P model and the GPT3.5 NoFineTune+Basic.P model, we found that the basic usage of ChatGPT (GPT-3.5-turbo with no fine-tuning and a basic prompt) may not achieve a desirable performance.
We find that our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) is able to generate test cases that are 78.5% being syntactically correct, 67.09% being aligned with the requirements while achieving a code coverage of 61.7%, e, and 18.9% mutation score which substantially outperforms all other LLMs). These results highlight the significant advancement of our approach for the text-to-testcase generation task when compared to other baseline LLMs.
Our approach GPT3.5FineTune+Imprv.P model emerges as the strongest model across all projects and evaluation metrics, particularly excelling in syntax correctness, requirement alignment, code coverage, and mutation scores. Gemini, while the next best performer, shows commendable results that are 72.35% syntactically correct, 62.94% aligned with the requirements while achieving a code coverage of 54.76% and a mutation score of 15.3%. Additionally, GPT-3.5+FineTunetUniTest incurs a cost for generating test cases. Conversely, for projects requiring more cost-effective solutions, Gemini presents a viable alternative that still maintains a comparable level of performance across all key metrics. The other models CodeGemma and CodeLlama are also consistently strong performers, particularly in FineTune+Imprv.P settings.
@@
When fine-tuned the LLMs (GPT-3.5-turbo) with our curated fine-tuned dataset, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT. With our improved prompts, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT.
When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see Figure 4). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 41.60% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
@@
{custombox}
Our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) is able to generate test cases that are 78.5% being syntactically correct, 67.09% being aligned with the requirements while achieving a code coverage of 61.7% and mutation score with 18.9%, which substantially outperforms all other LLMs.
@@
RQ2: What is the effect of the fine-tuning and prompting step for the LLM-based text-to-testcase generation task?
Figure 4: (RQ2)The experimental results of the ablation study with GPT3.5 model
@@
When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see figure 4). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 28.90% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of requirement alignment, we found that the performance is improved from 11.10% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 20.25% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of code coverage, we found that the performance is improved from 10.44% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 16.43% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 NoFine-Tune +Imprv.P, GPT3.5FineTune+Imprv.P and GPT3.5FineTune+Basic.P. Figure 4 illustrates that optimal results are achieved through the integration of fine-tuning and enhanced prompt engineering with GPT-3.5. It underscores the necessity of incorporating both fine-tuning and effective prompt design as essential strategies when deploying large language models for text-to-testcase generation tasks.
We employ the Wilcoxon Signed-Rank Test to evaluate the impact of fine-tuning and prompting on the performance metrics syntax correctness, requirement alignment, code coverage and mutation score(SC_Score, RA_Score, CC_Score, MS_Score). This non-parametric test is used to assess the differences in the median values. We use the test to estimate the relative contribution and the statistical significance (p-value). The tests compare the metrics under two conditions:
With and without fine-tuning comparison assesses whether the fine-tuning of the model significantly affects the performance metrics. With and without prompting comparison evaluates the effectiveness of prompting strategies in enhancing model performance. The results indicate statistically significant differences in all metrics for both conditions, with p-values significantly less than 0.05. This suggests that both fine-tuning and prompting contribute positively to improving model performance across various metrics.
Table 5: Wilcoxon Signed-Rank Test results for model performance.
Comparison Metric Statistic P-value
Prompting SC_Score 21.0 0.0021
RA_Score 25.0 0.0027
CC_Score 33.0 0.0133
MS_Score 24.0 0.0011
Fine-Tuning SC_Score 23.0 0.0012
RA_Score 27.0 0.0023
CC_Score 40.0 0.0136
MS_Score 40.0 0.0136
{custombox}
When fine-tuning the LLMs (GPT-3.5-turbo) with our curated fine-tuning data set, we observed an improvement of syntax correctness by 223%, the alignment of the requirements by 164%, the coverage of the code by 153%, and mutation score improved by 273% compared to the GPT-3.5-turbo model without the fine-tuning component. When considering our improved prompt design, we observed an improvement in syntax correctness by 124%, requirement alignment by 82%, code coverage by 58%, and mutation score improved by 150% compared to the GPT-3.5-turbo model without the fine-tuning component.
@@
RQ3: What types of errors occur in the incorrectly generated test cases?
Results. The majority of the generated test cases by our approach are syntactically correct for 78.5%, while being incorrect for 21.5%. The most common types of errors are assertion errors (11.3%), value errors (2.4%), syntax errors (6.9%), and others (0.9%).
@@
Understanding the incorrectly generated test cases by our approach is crucial as it helps researchers gain a better understanding of the limitations of our approach. To do so, we utilized Apache Maven [59] as our build tool, and the Surefire Plugin [60] to execute the test cases and to analyze the types of errors during the build process. Figure 5 provides a statistical summary of the error types of the generated test cases. An example of different types of incorrectly generated test cases is also presented in Listing 1. Below, we discuss the different types of errors.
[…]
  • The other (0.9%) category includes exception errors and runtime errors. The exception errors occur when the test case throws an unexpected exception or fails to handle an expected exception, such as NullPointerException, IndexOutOfBoundsException, IllegalArgumentException. The runtime errors also occur during the execution of the test case, such as NullPointerException or IllegalStateException.
@@
// Other error - Try to access value, which is null, leading to a NullPointerException
@Test
public void testJsonNullConstructor() {
JsonNull jsonNull = new JsonNull();
assertNotNull(jsonNull);
String value = jsonNull.getValue();
assertEquals("null", value);}
@@
{custombox}
found that a significant portion of the test cases generated by our approach exhibited syntactical correctness, with 78.5% of the generated test cases are syntactically correct, while 21.5% of the generated test cases are incorrect. Among the 21. 5% of incorrectly generated test cases, we found that the errors are related to assertion errors for 11.3%, value errors for 2.4%, syntax errors for 6.9%, and others for 0.9%.
6 Discussion
6.1 Recommendations to practitioners.
Consider Alice, a software developer engaged in a project that embraces Test-Driven Development (TDD) practices. In TDD, developers first write a failing test that defines a desired improvement or new function, then produce code to pass the test, and finally refactor the new code to acceptable standards. While this process inherently boosts software quality and aligns development with user requirements, it is also labor-intensive and demands high accuracy in test creation. Unfortunately, most automated test case generation tools available are designed primarily to work with existing code rather than generating tests from scratch based on new requirements. This presents a significant challenge and points to the need for tools capable of generating test cases directly from textual requirements descriptions. To bridge this gap, practitioners are encouraged to employ Large Language Models (LLMs) specifically fine-tuned for Text-to-Testcase generation tasks. The effectiveness of such models is notably enhanced when they are fine-tuned and utilized with well-crafted prompts. The results of RQ2 shows that the basic GPT-3.5 models, when fine-tuned and used with improved prompts, significantly outperform those that are not fine-tuned and employ basic prompts. By integrating these advanced LLMs into their workflow, developers can enhance their efficiency in test case generation, thereby adhering more effectively to TDD practices.
6.2 Usability and practicality.
To understand the usability and acceptability of the test cases generated by our approach to text-to-testcase generation, we conducted a user study. We aim to unveil the usability and acceptability of the test cases generated by our approach. We conduct our user study according to the following steps: (1) design and develop a survey, (2) select participants, and (3) survey questions. We explain the details of each step below.
(i) Design and Development of the Survey: Our survey utilises a cross-sectional design, capturing responses from participants at a singular point in time. Comprising four closed-ended and one open-ended question, the survey is structured to be completed anonymously in approximately 20 minutes. The closed-ended questions use a Likert scale format the scale ranges from 1 (strongly disagree) to 5 (strongly ggree), where each number corresponds to a specific level of agreement or satisfaction, and the survey itself is divided into two sections: initial demographic questions and questions regarding the usability and acceptability of test cases generated by our approach. We employed Qualtrics as the platform for our online survey administration, participants encountered a detailed introductory statement upon entering the survey, which outlines the study’s goals, the basis for participant selection, potential benefits and risks, and our dedication to maintaining confidentiality. Importantly, our survey underwent a rigorous evaluation process and received ethical approval from the Monash University Human Research Ethics Committee (MUHREC ID: 45001).
During the survey, participants were presented with test cases generated by the fine-tuned GPT3.5 our approach. For the five open-source projects we randomly selected two test cases from each project, resulting in a total of 10 test cases. Participants were asked to evaluate the usability and acceptability of each test case by answering four questions: the first focused on determining which test case was easier to read; the second aimed to identify relevance and correctness of the test case; the third asked participants to express their willingness to accept and adopt the generated test case for the given method description in one of their own projects, considering its relevance and quality; the fourth focus on the any other feedback or comments. (ii) Participant Selection: The target population for our study is Java software practitioners. We reached out to potential participants through LinkedIn and Facebook. Participants were invited to complete the online survey, which assessed the readability,usability and acceptability of test cases. Finally, we obtained a total of 61 responses in two weeks.
(iii) Survey Questions: The survey begins with a question about the participant’s role within their software development team (”(Q1). What is your role in the software development team?”). The next question concerns the years of experience the participants have in Java programming (”(Q2). How many years of experience do you have in Java programming?”), aimed at achieving a diverse range of responses from software practitioners with varied levels of professional experience.
The survey included precisely four questions presented to the participants: ”(Q1) How would you rate the readability of the generated test case?”; followed by (Q2) ”How would you rate the usability of the generated test cases?”; additionally (Q3) ”How would you to rank the acceptability of the generated test cases?”; finally any other feedback or comments.
To ensure the completeness of survey responses, especially for open-ended questions, we conducted a detailed manual review. Out of a total of 61 responses, we identified 8 invalid responses—due to unanswered open-ended questions or incomprehensible replies—and excluded them from further analysis. Consequently, we included the remaining 53 valid responses for analysis. The closed-ended responses were quantitatively analyzed.
The figure 6 presents the results of a survey using a 5-point Likert scale to assess the readability, usability, and acceptability of the generated test cases. The Likert scale ranges from 1 (Strongly Disagree or Very Poor) to 5 (Strongly Agree or Excellent). No respondents were rated Very Poor or Poor (ratings 1 and 2), indicating that there was no negative feedback. A minority of the respondents were neutral (rating 3), with percentages ranging from 14% to 16%. A larger proportion agreed that the test cases were good (rating 4), with scores ranging from about 26% to 30%. The majority felt strongly positive, rating the test cases as Excellent (rating 5), with significant majority above 50%. This distribution suggests a general positive reception for the usability and acceptability of text-to-testcase generation.
Figure 6: Survey Results of usability and acceptability of the test cases generated.
@@ 7 Threats to Validity
In our fine-tuning step, we employed a variety of hyperparameter configurations, including adjustments in the number of epochs, batch size, and learning rate. The hyperparameter optimization can significantly improve evaluation outcomes [61].
However, identifying the optimal settings is resource-intensive due to the expansive search space of hyperparameters. We note that our primary objective in this study is not to pinpoint the most effective hyperparameters but to investigate the potential of LLMs in TDD. Thus, while the results achieved with LLMs in our study are promising, they only represent a starting point and have the potential for further improvement through extensive hyperparameter tuning. Nonetheless, we have included detailed information about the hyperparameter settings used in our study within our replication package. The second major concern related to LLMs is the training data leakage [62]. There is a possibility that GPT-3.5-turbo or other LLMs may have been exposed to the same open-source projects that we used in our study. However, the pre-trained models did not yield satisfactory results, compared to the fine-tuned model and custom prompt results. This shows that the impact of training data leakage, if any, was minimal in our study settings.
Construct Validity. Our study’s main construct validity concern is related to the prompt tuning and robustness [63]. Given the large scope of subjectivity in prompting and the lack of empirical guidelines on prompt tuning and robustness in SE, one can always further tune and build a more robust prompt to achieve better evaluation results. However, we experimented with several prompt combinations in our journey, from the basic to the custom prompt, in an attempt to mitigate the threats related to prompt tuning. The second major concern is related to the potential data leakage (i.e., when instances in the training dataset appear in the evaluation set) of our model. To mitigate this threat, we focus on identifying instances in the training dataset (i.e., method2test) that appear in the evaluation set (i.e., open-source projects). For the method2test dataset, we discovered and removed 63 instances for the JFreeChart project. This will ensure that there is no data leakage in our experiments.